# Chinese Academy of Sciences, Beijing, China

### v1

@@ 2 Preliminaries
$f(2^{\omega}):=\{\hskip 0.85355pt{y}:{\exists x,\ h(x)=y}\hskip 0.85355pt\}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf^{-1}(y):=\{\hskip 0.85355pt{x}:{f(x)=y}\hskip 0.85355pt\}$
@@ 3 Deterministic inversions
We define a computable representation $\hat{g}$ of the required $g$. Let
$B_{\tau}^{n}:=\{\hskip 0.85355pt{\rho\in 2^{n}}:{\hat{f}(\rho)\preceq\tau}\hskip 0.85355pt\}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptp_{\tau}:=\max\{\hskip 0.85355pt{n}:{|{B_{\tau}^{n}}|=1}\hskip 0.85355pt\}$
and define $\hat{g}(\tau)$ to be the unique string in $B^{p_{\tau}}_{\tau}$. Then $p_{\tau}$ is nondecreasing, $\hat{g}$ is $\preceq$-monotone and by (1) both $p_{\tau},\hat{g}$ are computable. So $\hat{g}$ is a computable representation of a partial computable $g:\subseteq 2^{\omega}\to 2^{\omega}$.
@@ 4 Randomized inversions
Since $p$ is computable the sets $f^{-1}(\llbracket V_{i}\rrbracket)$ are uniformly $\Sigma^{0}_{1}$. Also
@@ 4.1 Construction of a one-way function
$t\geq k\ \Rightarrow\ \frac{\mu(\llbracket(y\oplus r)\upharpoonright_{2t})\rrbracket)\cap L)}{\mu(\llbracket(y\oplus r)\upharpoonright_{2t})\rrbracket)}>\frac{1}{2}.$ (3)
@@ 4.2 Injectivity and one-way functions
We consider the extent to which one-way functions fail to be injective. Our observations hold more generally with respect to inversions on one variable.
@@
Every (weakly) one-way $f$ is almost nowhere injective.
@@
If $f:2^{\omega}\to 2^{\omega}$ is (weakly) one-way then $f^{-1}(y)$ is uncountable for almost every $y\in f(2^{\omega})$.

### v2

@@ 2 Preliminaries
$f(2^{\omega}):=\{\hskip 0.85355pt{y}:{\exists x,\ f(x)=y}\hskip 0.85355pt\}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf^{-1}(y):=\{\hskip 0.85355pt{x}:{f(x)=y}\hskip 0.85355pt\}$
@@ 3 Deterministic inversions
Let $A_{\tau}=\{\sigma:\hat{f}(\sigma)\succeq\tau\}$, $B_{\tau}$ be the $\preceq$-downward closure of $A$,
$B_{\tau}^{n}:=B_{\tau}\cap 2^{n}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptp_{\tau}:=\max\{\hskip 0.85355pt{n}:{|{B_{\tau}^{n}}|=1}\hskip 0.85355pt\}.$
We define a computable representation $\hat{g}$ of the required $g$ by letting $\hat{g}(\tau)$ to be the unique string in $B^{p_{\tau}}_{\tau}$. Then $p_{\tau}$ is nondecreasing, $\hat{g}$ is $\preceq$-monotone and by (1) $B_{\tau}^{n}$, $p_{\tau}$ and $\hat{g}$ are computable. So $\hat{g}$ is a computable representation of a partial computable $g:\subseteq 2^{\omega}\to 2^{\omega}$.
@@ 4 Randomized inversions
Since $f$ is computable the sets $f^{-1}(\llbracket V_{i}\rrbracket)$ are uniformly $\Sigma^{0}_{1}$. Also
@@ 4.1 Construction of a one-way function
$t\geq k\ \Rightarrow\ \frac{\mu(\llbracket(y\oplus r)\upharpoonright_{2t})\rrbracket)\cap L)}{\mu(\llbracket(y\oplus r)\upharpoonright_{2t})\rrbracket)}>\frac{1}{2}$ (3)
@@ 4.2 Injectivity and one-way functions
We consider the extent to which one-way functions fail to be injective. Our observations hold generally with respect to inversions on one variable.
@@
Every (weakly) one-way total computable function is almost nowhere injective.
@@
If $f:2^{\omega}\to 2^{\omega}$ is a total computable (weakly) one-way function then $f^{-1}(y)$ is uncountable for almost every $y\in f(2^{\omega})$.